OpenStudy (anonymous):

Choice A: 3 nights and one meal for $ 250 Choice B: 3 nights and 6 meals for $330 A motel offers the specials shown. What price is the motel charging per night?

OpenStudy (anonymous):

3n + m = 250 -- (3 nights(n) + 1 meal (m) = 250) --- multiply by -1 3n + 6m = 330 -- (3 nights + 6 meals = 330) --------------------- -3n -m = -250 (result of multiplying by -1) 3n + 6m = 330 --------------------add 5m = 80 m = 80/5 m = 16 now sub 16 in for n in either of the original equations 3n + m = 250 3n + 16 = 250 3n = 250 - 16 3n = 234 n = 234/3 n = 78 check.. 3n + 6m = 330 3(78) + 6(16) = 330 234 + 96 = 330 330 = 330 (correct) so each night (n) costs $ 78